# Stable Symmetry-Protected Topological Phases in Systems with Heralded Noise

### v1

@@ V Mean field analysis
$n^{d}=\begin{dcases}\ \frac{1-f_{e}}{2}+\frac{\eta f_{e}^{2}}{4\gamma}\qquad&\text{if }\ \eta f_{e}<2\gamma,\\ \ 1/2\qquad&\text{if }\ \eta f_{e}\geq 2\gamma\end{dcases}$ (S37)
@@ VI.1 Number of erasures
$\begin{split}\mathcal{L}_{\text{contact}}(\rho^{e})=\eta&\sum_{j}\left(e_{j}^{-}\rho^{e}e_{j}^{+}-\frac{1}{2}\{1-n^{e}_{j},\rho^{e}\}\right)+\gamma\sum_{j}\left(e^{+}_{j-1}(1-n^{e}_{j+1})\rho^{e}(1-n^{e}_{j+1})e^{-}_{j-1}-\frac{1}{2}\{n^{e}_{j-1}(1-n^{e}_{j+1}),\rho^{e}\}\right)\\ &+\gamma\sum_{j}\left(e^{+}_{j+1}(1-n^{e}_{j-1})\rho^{e}(1-n^{e}_{j-1})e^{-}_{j+1}-\frac{1}{2}\{n^{e}_{j+1}(1-n^{e}_{j-1}),\rho^{e}\}\right).\end{split}$ (S43)
@@
Figure S3: Critical exponents for the stabilizers defects: (a) The deviation of the density of stabilizer defects $n_{d}$ from its value in the absorbing state as a function of time. The data is presented for $L=512$ and varying noise rates close to the critical point. The critical curve shows a power law decay $\sim t^{-\delta_{d}}$. The estimation of exponents $\delta_{d}$ and $\nu_{t}$ is shown in Fig. 2 in the main text. (b) The dynamical exponent $z=1.58(4)$ is estimated by observing the data collapse of the critical curve $\eta=\eta_{c}$ for varying system sizes. The data is obtained by averaging over $2\times 10^{4}$ independent Monte Carlo runs initialized in the perfectly ordered state $S_{j}=1,n^{e}_{j}=0$ for all sites $j$.
@@ VI.2 Density of stabilizer defects
$n^{d}(t):=\text{Tr}{\left(\rho(t)\frac{1}{L}\sum_{j=1}^{L}n^{d}_{j}\right)}=\frac{1}{L}\sum_{j=1}^{L}\text{Tr}{\left(\rho(t)\frac{1-S_{2j}}{2}\right)}.$ (S44)
@@ VI.3 String-order
Figure S4: String order critical exponents: (a) the average value of the string order for length $L/2$ (see Eq. (S46)) is plotted as a function of time for noise rates close to critical point. The string order at the critical noise rate decays with a power law $t^{-0.32}$, consistent with the exponent $\delta_{\Omega}$ obtained in the main text. The data is plotted for $L=512$ sites per sublattice and $2\times 10^{4}$ independent Monte Carlo runs. (b) The finite size collapse at the critical noise-rate $\eta_{c}$ is observed when the dynamical exponent is tuned to $z=1.58(9)$. (c) The steady state value of the string order $\Omega_{l}$ is plotted as a function of the length of the string $l$ for an $L=512$ sized system. The data for a wide range of noise rates matches with the steady state value of the two point correlation function of erasures defined by $\langle(1-n^{e}_{1})(1-n^{e}_{l})\rangle$ (shown using dashed lines of the same colors).
@@ VII Initial state dependence
$c_{0}=\frac{1}{\text{Tr}{(l_{0}^{\dagger}r_{0})}}\text{Tr}{\left(\hat{l}^{\dagger}_{0}\rho_{0}\right)}=1,\qquad c_{1}=\frac{1}{\text{Tr}{(l_{1}^{\dagger}r_{1})}}\text{Tr}{\left(\hat{l}^{\dagger}_{1}\rho_{0}\right)}\propto\text{Tr}{\left(\tau^{e}_{L}\rho_{0}\right)\,.}$ (S54)
@@ VIII.1 Dynamics of Unheralded Defects
$\frac{d}{dt}h_{\text{noise}}=2\eta(1-f_{e})(1-n^{e})^{2}\ -\ 2\eta(2-f_{e})h.$ (S61)
@@
In particular, let us focus on the parameter regime $\eta\langle\langle\gamma$ and $f_{e}\sim 1$. The erasure noise channel generates the configuration $b$ by adding a flag next to an isolated unheralded defect. This channel can also reduce the density of $b$ by adding flags and converting unheralded defects into heralded defects. The resulting dynamics are approximated as
@@
Note that higher order correction processes can also lead to the creation of $b$ configurations. One such process corresponds to the shortening of a longer erasure string with only one defect at its endpoint. Here, we neglect such terms since they give sub-leading contributions when $\eta\langle\langle\gamma$. Combining the contributions from Eqs. (S63), (S64), and (S65), we obtain the steady state density of $b$ as:
@@
where we have assumed that $n^{e}\langle\langle 1$ for parameters that are well inside the ordered phase ($\eta\langle\langle\gamma,f_{e}\sim 1$). The numerical values of the steady state density of unheralded defects obtained using the Monte Carlo simulations for $\eta/\gamma=0.3$ are shown in Fig. S8a. The density grows $\propto\sqrt{1-f_{e}}$, as predicted by the mean-field equation that only considers dominant higher order processes.
Figure S9: Finite time signature of criticality: (a) The string order is plotted as a function of time for parameters close to the critical point $\eta/\gamma=\eta_{c}$ and for various values of the heralding fraction $f_{e}$. The datapoints are obtained using Monte Carlo simulations for an $L=256$ sized system initialized in the cluster state. The dashed line show the best-fits to $at^{-\delta_{\Omega}}e^{-bt}$ where $a$ and $b$ are fit parameters and $\delta_{\Omega}=0.32$ is the order-parameter decay exponent. (b) The fit parameter $a$ takes a constant value close to the pre-factor obtained for the $f_{e}=1$ case (0.55). The parameter $b$ matches well with $4\eta(1-f_{e})$ which corresponds to the rate of decay of $\Omega$ in the presence of unheralded $Z$ noise-channel with a rate $\eta(1-f_{e})$. This captures the cross-over from an initial critical power law decay $t^{-\delta_{\Omega}}$ to a finite-rate exponential decay $e^{-4\eta(1-f_{e})t}$ that is dominated by the presence of unheralded defects.
Now, we estimate the steady state value of the string order in subregion of length $l$. When the noise rate is low $\eta\langle\langle\gamma$ and the fraction of unheralded errors is small $(f_{e}\sim 1)$, we expect that there will be two contributions to the density of stabilizer defects. The defects that are confined due to the flags will give a finite contribution to the string order that is independent of the string length $l$. In contrast, stabilizer defects that are not heralded by flags (previously denoted by $h$) will remain deconfined. These uncorrelated defects destroy the string order of longer strings in the steady state. We estimate the string order $\Omega_{l}$ in terms of the unheralded defect density as
@@
where we have approximated $\log{(1-2h)}\sim-2h$ for a small value of the unheralded defect density. Using this result, we estimate the string-order length-scale $\xi_{\Omega}\propto 1/h\propto(1-f_{e})^{-1/2}$. The estimated variation of $\xi_{\Omega}$ as a function of $f_{e}$ matches well with the numerical results reported in Fig. S8c and Fig. 3c in the main text. Finally, in Figs. S9a and S9b we show the finite-time behavior of the string order, which displays a cross-over from an initial power law decay (dictated by $\delta_{\Omega}$, the order-parameter decay exponent) to an exponential decay (dictated by unheralded errors), with the cross-over occurring at a time-scale $\sim\frac{1}{\eta(1-f_{e})}$.
@@ IX Alternative approach to stabilizing string-order
The chiral dynamics pushes the stabilizer defects towards the center of the system, bringing them close together and allowing for local correction. Let us first focus on sites in the bulk of the system that are away from the center and from the endpoints. The exact time evolution of the density of defects at such sites $k$ is governed by
@@
Figure S10: Stabilization of string order using biased motion: (a) the string order parameter $\Omega_{2,L-2}:=X_{3}Z_{4}Z_{6}\ldots Z_{2L-4}X_{2L-3}$ is plotted as a function of time for varying values of the drift parameter $\mu$. The data is shown for $L=256$ sites per sub-lattice and noise rate $\eta/\gamma=0.1$. (b) The steady state value of the string order parameter as a function of the location of the left end-point $k$. The midpoint of the string is at the center of the system. The dashed lines show the best fit to an exponential function $e^{-(k-1)/\xi}$. The data-points in the inset show the length-scale $\xi$ obtained from the fits as a function of $\mu$ for different values of $\eta$. The values match well with the analytical prediction in Eq. (S83) shown by the dashed curve. (c) The average number of stabilizer defects in the steady state at each site is plotted as function of the site-location. The number of defects per site in the bulk approaches the predicted value (Eq. (S77)) shown using the magenta line. The dotted lines joining the numerical data points are a guide to the eye. The variation of the number of defects at the boundary sites is plotted in the inset as a function of $\mu$. The data-points match well with the predicted curve (Eq. (S79)), shown by dashed lines of same color. The numerical results are obtained by Monte Carlo simulations of the Lindblad master equation in the population basis of the stabilizers on even sites (the approximations for the heralded noise model discussed in Sec. IV are valid for this model as well). The final results in the plots are obtained by averaging over $10^{4}$ independent Monte Carlo samples initialized in the cluster state.
@@
where $a$ is the spacing between sites and where the additional factor of one-half indicates that the length of the string is only modified from one side. The first set of equalities in the equation above result from the invariance of the dynamics with respect to the reflection symmetry about the center of the chain. Using this continuum approximation, along with Eq. (S80), we obtain the steady state equation for the string order as
@@
Notice that $\Omega(x=0)$ can be written as $\prod_{i=1}^{L-1}S_{2i}$. As a result of the dynamics that does not allow stabilizer defects to leak through the boundary (Eq. (S74)), this operator is conserved and takes the value $1$ if the system is initialized in the cluster state. Hence, by definition $\Omega$ across the entire string is equal to $1$ for all times. Using this boundary condition, we get the steady state string order to be
@@
This implies that there is a finite steady-state string order when both the left and right endpoints of the string are within a distance $\xi$ of the respective system-boundaries. Since the length of such a string is equal to $L-2x$, we have a steady state that has a boundary sensitive long-range string order. In Fig. (S10a), we numerically verify the long string order for a system initialized in the cluster state. Additionally, the analytical mean-field estimate for the length scale $\xi$ also matches well with the numerical data for a wide range of parameters (see Fig. (S10b)).

### v3

@@ V Mean field analysis
$n^{d}=\begin{dcases}\ \frac{1-f_{e}}{2-f_{e}}+\frac{\eta f_{e}^{2}}{4\gamma(2-f_{e})}\qquad&\text{if }\ \eta f_{e}<2\gamma,\\ \ 1/2\qquad&\text{if }\ \eta f_{e}\geq 2\gamma\end{dcases}$ (S37)
@@ VI.1 Number of erasures
$\begin{split}\mathcal{L}_{\text{contact}}(\rho^{e})=\eta&\sum_{j}\left(e_{j}^{-}\rho^{e}e_{j}^{+}+n^{e}_{j}\rho^{e}n^{e}_{j}-\rho^{e}\right)+\gamma\sum_{j}\left(e^{+}_{j-1}(1-n^{e}_{j+1})\rho^{e}(1-n^{e}_{j+1})e^{-}_{j-1}-\frac{1}{2}\{n^{e}_{j-1}(1-n^{e}_{j+1}),\rho^{e}\}\right)\\ &+\gamma\sum_{j}\left(e^{+}_{j+1}(1-n^{e}_{j-1})\rho^{e}(1-n^{e}_{j-1})e^{-}_{j+1}-\frac{1}{2}\{n^{e}_{j+1}(1-n^{e}_{j-1}),\rho^{e}\}\right).\end{split}$ (S43)
@@
Figure S4: Critical exponents for the stabilizers defects: (a) The deviation of the density of stabilizer defects $n^{d}$ from its value in the absorbing state as a function of time. The data is presented for $L=512$ and varying noise rates close to the critical point. The critical curve shows a power law decay $\sim t^{-\delta_{d}}$. The estimation of exponents $\delta_{d}$ and $\nu_{t}$ is shown in Fig. 2 in the main text. (b) The dynamical exponent $z=1.58(4)$ is estimated by observing the data collapse of the critical curve $\eta=\eta_{c}$ for varying system sizes. The data is obtained by averaging over $2\times 10^{4}$ independent Monte Carlo runs initialized in the perfectly ordered state $S_{j}=1,n^{e}_{j}=0$ for all sites $j$.
@@ VI.2 Density of stabilizer defects
$n^{d}(t):=\text{Tr}{\left(\rho(t)\frac{1}{L}\sum_{j=1}^{L}n^{d}_{2j}\right)}=\frac{1}{L}\sum_{j=1}^{L}\text{Tr}{\left(\rho(t)\frac{1-S_{2j}}{2}\right)}.$ (S44)
@@ VI.3 String-order
Figure S5: String order critical exponents: (a) the average value of the string order for length $L/2$ (see Eq. (S46)) is plotted as a function of time for noise rates close to critical point. The string order at the critical noise rate decays with a power law $t^{-0.32}$, consistent with the exponent $\delta_{\Omega}$ obtained in the main text. The data is plotted for $L=512$ sites per sublattice and $4\times 10^{4}$ independent Monte Carlo runs. (b) The finite size collapse at the critical noise-rate $\eta_{c}$ is observed when the dynamical exponent is tuned to $z=1.58(9)$. The data is obtained by averaging $2\times 10^{4}$ Monte Carlo realizations. (c) The steady state value of the string order $\Omega_{l}$ is plotted as a function of the length of the string $l$ for an $L=512$ sized system. The data for a wide range of noise rates matches with the steady state value of the two point correlation function of erasures defined by $\langle(1-n^{e}_{1})(1-n^{e}_{l})\rangle$ (shown using dashed lines of the same colors). The steady state values are obtained by simulating $5\times 10^{3}$ realizations upto $2L^{1.5}$ MC sweeps.
@@ VII Initial state dependence
$c_{0}=\text{Tr}{\left(\hat{l}^{\dagger}_{0}\rho_{0}\right)}=1,\qquad c_{1}=\text{Tr}{\left(\hat{l}^{\dagger}_{1}\rho_{0}\right)}\propto\text{Tr}{\left(\tau^{e}_{L}\rho_{0}\right)\,.}$ (S54)
@@ VIII.1 Dynamics of Unheralded Defects
$\frac{d}{dt}h\bigg{|}_{\text{noise}}=2\eta(1-f_{e})(1-n^{e})^{2}\ -\ 2\eta(2-f_{e})h.$ (S61)
@@
In particular, let us focus on the parameter regime $\eta\ll\gamma$ and $f_{e}\sim 1$. The erasure noise channel generates the configuration $b$ by adding a flag next to an isolated unheralded defect. This channel can also reduce the density of $b$ by adding flags and converting unheralded defects into heralded defects. The resulting dynamics are approximated as
@@
Note that higher order correction processes can also lead to the creation of $b$ configurations. One such process corresponds to the shortening of a longer erasure string with only one defect at its endpoint. Here, we neglect such terms since they give sub-leading contributions when $\eta\ll\gamma$. Combining the contributions from Eqs. (S63), (S64), and (S65), we obtain the steady state density of $b$ as:
@@
where we have assumed that $n^{e}\ll 1$ for parameters that are well inside the ordered phase ($\eta\ll\gamma,f_{e}\sim 1$). The numerical values of the steady state density of unheralded defects obtained using the Monte Carlo simulations for $\eta/\gamma=0.3$ are shown in Fig. S9a. The density grows $\propto\sqrt{1-f_{e}}$, as predicted by the mean-field equation that only considers dominant higher order processes.
Figure S10: Finite time signature of criticality: (a) The string order is plotted as a function of time for parameters close to the critical point: $\eta/\gamma=\eta_{c}$ and for various values of the heralding fraction $f_{e}$. The datapoints are obtained using Monte Carlo simulations for an $L=256$ sized system initialized in the cluster state. The dashed line show the best-fits to $at^{-\delta_{\Omega}}e^{-bt}$ where $a$ and $b$ are fit parameters and $\delta_{\Omega}=0.32$ is the order-parameter decay exponent. (b) The fit parameter $a$ takes a constant value close to the pre-factor obtained for the $f_{e}=1$ case (0.55). The parameter $b$ matches well with $4\eta(1-f_{e})$ which corresponds to the rate of decay of $\Omega$ in the presence of unheralded $Z$ noise-channel with a rate $\eta(1-f_{e})$. This captures the cross-over from an initial critical power law decay $t^{-\delta_{\Omega}}$ to a finite-rate exponential decay $e^{-4\eta(1-f_{e})t}$ that is dominated by the presence of unheralded defects.
Now, we estimate the steady state value of the string order in subregion of length $l$. When the noise rate is low $\eta\ll\gamma$ and the fraction of unheralded errors is small $(f_{e}\sim 1)$, we expect that there will be two contributions to the density of stabilizer defects. The defects that are confined due to the flags will give a finite contribution to the string order that is independent of the string length $l$. In contrast, stabilizer defects that are not heralded by flags (previously denoted by $h$) will remain deconfined. These uncorrelated defects destroy the string order of longer strings in the steady state. We estimate the string order $\Omega_{l}$ in terms of the unheralded defect density as
@@
where we have approximated $\log{(1-2h)}\sim-2h$ for a small value of the unheralded defect density. Using this result, we estimate the string-order length-scale $\xi_{\Omega}\propto 1/h\propto(1-f_{e})^{-1/2}$. The estimated variation of $\xi_{\Omega}$ as a function of $f_{e}$ matches well with the numerical results reported in Fig. S9c and Fig. 3c in the main text. Finally, in Figs. S10a and S10b we show the finite-time behavior of the string order, which displays a cross-over from an initial power law decay (dictated by $\delta_{\Omega}$, the order-parameter decay exponent) to an exponential decay (dictated by unheralded errors), with the cross-over occurring at a time-scale $\sim\frac{1}{\eta(1-f_{e})}$.
@@ IX Alternative approach to stabilizing string-order
The chiral dynamics pushes the stabilizer defects towards the center of the system, bringing them close together and allowing for local correction. Let us first focus on sites in the bulk of the system that are away from the center and from the endpoints (consider left half of the chain, calculations work similarly for right half due to reflection symmetry). The exact time evolution of the density of defects at such sites $k$ is governed by
@@
Figure S11: Stabilization of string order using biased motion: (a) the string order parameter $\Omega_{2,L-2}:=X_{3}Z_{4}Z_{6}\ldots Z_{2L-4}X_{2L-3}$ is plotted as a function of time for varying values of the drift parameter $\mu$. The data is shown for $L=256$ sites per sub-lattice and noise rate $\eta/\gamma=0.1$. (b) The steady state value of the string order parameter as a function of the location of the left end-point $k$. The midpoint of the string is at the center of the system. The dashed lines show the best fit to an exponential function $e^{-(k-1)/\xi}$. The data-points in the inset show the length-scale $\xi$ obtained from the fits as a function of $\mu$ for different values of $\eta$. The values match well with the analytical prediction in Eq. (S83) shown by the dashed curve. (c) The average number of stabilizer defects in the steady state at each site is plotted as function of the site-location. The number of defects per site in the bulk approaches the predicted value (Eq. (S77)) shown using the magenta line. The dotted lines joining the numerical data points are a guide to the eye. The variation of the number of defects at the boundary sites is plotted in the inset as a function of $\mu$. The data-points match well with the predicted curve (Eq. (S79)), shown by dashed lines of same color. The numerical results are obtained by Monte Carlo simulations of the Lindblad master equation in the population basis of the stabilizers on even sites (the approximations for the heralded noise model discussed in Sec. IV are valid for this model as well). The steady state results in the plots are evaluated after $4\times 10^{3}$ MC sweeps by averaging over $10^{4}$ independent Monte Carlo samples initialized in the cluster state.
@@
where $a$ is the size of unit cell, and where the additional factor of one-half indicates that the length of the string is only modified from one side. The first set of equalities in the equation above result from the invariance of the dynamics with respect to the reflection symmetry about the center of the chain. Using this continuum approximation, along with Eq. (S80), we obtain the steady state equation for the string order as
@@
Notice that $\Omega(x=0)$ can be written as $\prod_{i=1}^{L-1}S_{2i}$. As a result of the dynamics that does not allow stabilizer defects to leak through the boundary (Eq. (S74)), this operator is conserved and takes the value $1$ if the system is initialized in the cluster state. Hence, by definition $\Omega$ across the entire string is equal to $1$ for all times. Using this boundary condition, we get the steady state string order (setting $a=1$) to be
@@
This implies that there is a finite steady-state string order when both the left and right endpoints of the string are within a distance $\xi$ of the respective system-boundaries. Since the length of such a string is equal to $L-2x$, we have a steady state that has a boundary sensitive long-range string order. In Fig. (S11a), we numerically verify the long-time string order for a system initialized in the cluster state. Additionally, the analytical mean-field estimate for the length scale $\xi$ also matches well with the numerical data for a wide range of parameters (see Fig. (S11b)).